OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work.

OpenStudy (bahrom7893):

standard form is: y-y1 = M(x-x1) So find M first.

OpenStudy (bahrom7893):

M = (y-y1)/(x-x1)

OpenStudy (anonymous):

m=-1?

OpenStudy (ja1):

*facedesk* @bahrom7893

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

You got the slope right. Now pick any of the two points, and use them as your (x1, y1) in this formula: y-y1 = M(x-x1)

OpenStudy (anonymous):

I got y=-x-3. Is that correct? @bahrom7893

OpenStudy (bahrom7893):

Which point did you use?

OpenStudy (bahrom7893):

If I use the second point: y-(-1) = -1(x-(-2)) y+1 = -(x+2) y+1 = -x-2 y = -x-3

OpenStudy (bahrom7893):

Correct. Any point you pick will give you the same answer btw.

OpenStudy (anonymous):

Okay Thank you